OpenStudy (anonymous):

log_{1/3} (log_{4} (x^2-5)>0

OpenStudy (anonymous):

@satellite73 u answer this yesterday can u explain the diff bet toays and yester

OpenStudy (anonymous):

todays has one log this one has two

OpenStudy (anonymous):

so ..........

OpenStudy (anonymous):

so we have to do it in two pieces

OpenStudy (anonymous):

if u can plzz can u do it again..... if u r not intrested plzz inform me

OpenStudy (anonymous):

can you find the question you asked last time? it should be in your profile

OpenStudy (anonymous):

let me check it out and will inform u

OpenStudy (anonymous):

we can do it again if you like first of all don't forget that the domain of any logarithmic function is \((0,\infty)\) i.e. you cannot take the log of a non- positive number so one thing we have to make sure of is that \(\log_4(x^2-5)>0\)

OpenStudy (anonymous):

that has nothing to do with the inequality that you have to solve, that has to do with the domain of the function we can solve that easily \(\log_4(x^2-5)>0\iff x^2-5>1\iff x^2-6>0\iff x<-\sqrt{6}\) or \(x>\sqrt{6}\)

OpenStudy (anonymous):

so we have that restriction before we even start to solve the inequality you are given, namely \[\log_{1/3} (\log_{4} (x^2-5))>0\]

OpenStudy (anonymous):

@satellite73 plz look my new ques http://openstudy.com/study?signup#/updates/4fbe5c0be4b0c25bf8fc84dd

OpenStudy (anonymous):

now we also know that if \(\log_{\frac{1}{3}}(x)>0\) we must have \(x<1\) so we also need to solve \[\log_4(x^2-5)<1\] \[x^2-5<4\] \[x^2-9<0\] \[-3<x<3\] put these together to get \[\sqrt{6}<x<3\]